A remote-controlled car is moving in a vacant parking lot. The velocity of the car as a function of time is given by 𝑣 = [5.00π‘š/𝑠 βˆ’ (0.0180π‘š/ 𝑠 3 )𝑑 2 ]𝑖̂+ [2.00 π‘šβ„π‘  + (0.550 π‘š 𝑠 2 ⁄ )𝑑]𝑗. (a) What are Μ‚ π‘Žπ‘₯(𝑑) and π‘Žπ‘¦(𝑑), the π‘₯- and 𝑦-components of the car’s velocity as functions of time? (b) What are the magnitude and direction of the car’s velocity at 𝑑 = 8.00 𝑠? (b) What are the magnitude and direction of the car’s acceleration at 𝑑 = 8.00 𝑠?


1
Expert's answer
2021-09-29T09:55:51-0400

(a) The x- and y-component of car's velocity:


"v_y=[2.00+ (0.550 )𝑡] \text{ m/s},\\
v_x=[5.00 − (0.0180)𝑡^2 ]\text{ m/s}."


Find the acceleration:


"a(t)=v'(t)=[-0.036t]\hat i+[0.55]\hat j.\\
a_x(t)=-0.036t,\\
a_y(t)=\text{const}
=0.55\text{ m/s}^2."

(b) The magnitude:


"v=\sqrt{(2+0.55·8)^2+(5-0.018·8^2)^2}=7.47\text{m/s},\\\space\\
\theta=\arctan\frac{2+0.55·8}{5-0.018·8^2}
=58.0°"

from +i toward +j.


"a=\sqrt{(-0.036·8)^2+(0.55)^2}=0.620\text{ m/s}^2.\\\space\\
\phi=\arctan\frac{0.55}{-0.036·8}=-62.4°"

from +i toward +j (or 62.4Β° from +i to +j).Β